OpenStudy (anonymous):

What is the solution of 3(2m-1) ≤ 4m+7? Please explain.

OpenStudy (anonymous):

no^^

OpenStudy (anonymous):

6m-3<=4x+7 -4m -4m 2m-3<=7 +3 +3 2m<=10 m<=5

OpenStudy (anonymous):

\[m \le5\]

OpenStudy (anonymous):

6m - 3 <= 4m + 7 2m <= 10 m <= 5

OpenStudy (anonymous):

Jimmyrep is right